Find the linearization of the function z=x√y at the point (-9, 25).
L(x,y)=
1
Expert's answer
2020-10-12T13:25:45-0400

f(x,y)=f(Z)= xy1/2

Fx of the function f(x,y) = y1/2

Fy of the function f(x,y) = (xy-1/2)/2

Now, linearization of a function f(x,y)=f(z) at (a,b) is given by

L(x,y) = f(a,b) + (x-a)fx(a,b) + (y-b)fy(a,b)

f(a,b) = f(-9,25)=(-9)(25)1/2= (-45)

fx(a,b) = fx(-9,25)=y1/2=(25)1/2=5

fy(a,b)=fy(-9,25)=(xy-1/2)/2=-9(25-1/2)/2

= -0.9

So, linearization of f(x,y) = xy1/2 at (a,b)

= (-9,25) is given by

L(x,y) = f(-9,25)+(x-(-9))fx(-9,25)+(y-25)

fy(-9,25)

L(x,y) = -45+(x+9)(5)+(y-25)(-0.9)

L(x,y) = -45+5x+45+(-0.9y+22.5)

L(x,y) = -45+5x+45-0.9y+22.5

L(x,y) = 5x-0.9y+22.5

so, linearization of f(x,y)=xy1/2 at (a,b)=

(-9,25) is given by

L(x,y) = 5x-0.9+22.5